A 918 kg car has four 12.7 kg wheels. When the car is moving, what fraction of the total kinetic energy of the car is due to rotation of the wheels about their axles? Assume that the wheels have the same rotational inertia as uniform disks of the same mass and size.

Grade:12th pass

1 Answers

Arun
5 years ago
in case the weight of wheels is included in the lump sum of 918 kg  
Ek of wheels = Ekt (translational)+Ekr (rotational) : 
Ek 4(1/2*m*V^2+1/2*J*ω^2) = 4(1/2m*V^2+1/2(1/2*m*r^2*V^2/r^2)) = 4*m*V^2(1/2+1/4) = 3*mV^2 = 38.1V^2 of which : 
# 38.1V^2/3 is rotational 
# 2*38.1V^2/3 is translational  

and finally : 
k1 = 100(38.1/(918/2+38.1/3)) = 
 
nownyou cqn calculate